# Deep reinforcement learning-based active flow control of an elliptical cylinder: transitioning from an elliptical cylinder to a circular cylinder and a flat plate

### v2

@@ II.2 Deep Reinforcement learning
In the domain of Reinforcement Learning, the interaction between an agent and its environment is the core mechanism by which learning is facilitated. As illustrated in Figure 3, this interaction proceeds in a cyclical pattern. At each timestep $t$, the agent observes the current state $s_{t}$ of the environment, which contains all the relevant information necessary for making decisions. Based on this observation, the agent selects and executes an action $a_{t}$, according to its policy $\pi$. This policy, which is essentially a strategy that maps states to actions, may function deterministically or probabilistically. After the action is executed, the environment transitions to a new state $s_{t+1}$ and awards the agent with an immediate reward $r_{t+1}$. This reward, quantified as a numerical value, evaluates the action’s efficacy in progressing towards the agent’s objectives. Subsequently, the agent assimilates this feedback—comprised of both the reward and the new state—to refine its policy, employing RL algorithms designed to optimize the accumulation of long-term rewards.
Deep Reinforcement learning
Deep learning, a specialized subset of machine learning, excels at identifying patterns in extensive datasets through its multi-layer neural network architecture. This approach enables computers to derive complex representations directly from raw data, leading to revolutionary advances in fields like image recognition, natural language processing, and predictive analytics. DRL marries the potent perceptual capabilities of deep learning with the decision-making and optimization capabilities of reinforcement learning, effectively overcoming their individual limitations. Deep learning addresses the challenge of processing original, high-dimensional input data, which is problematic for traditional reinforcement learning, while reinforcement learning provides a structured decision-making framework and a goal-oriented learning mechanism that enhances the capabilities of deep learning. As depicted in Figure 3, a deep learning network transforms the intricate state of the environment into a compressed and efficient feature representation. This transformation yields a more manageable and interpretable state space for reinforcement learning algorithms. In DRL, reinforcement learning leverages these high-level features to learn and optimize decision-making policies. These policies are typically parameterized by deep neural networks, implying that the selection of actions is influenced not only by the current observation of the environmental state but also by the learned weights within the network.
II.3 Proximal Policy Optimization algorithm
PPO is an advanced DRL algorithm suitable for addressing reinforcement learning challenges within both discrete and continuous action spaces. The fundamental concept of PPO revolves around enhancing performance by maximizing a surrogate objective function while ensuring the stability of policy updates. In PPO, an agent interacts with an environment to gather experience data, which is then utilized to construct the objective function. This function incorporates dominant estimates from the experience data to gauge the direction of policy improvement. During these interactions, the agent generates a sequence of trajectories ($\tau$). A trajectory consists of a series of state-action pairs produced as the agent operates under the current policy, continuing until a terminal state is reached or a predetermined maximum number of steps is exceeded. Symbolically, a trajectory $\tau$ can be represented as follows:
$\tau=(s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},s_{2},\ldots,s_{T-1},a_{T-1},r_{T},s_{T}),$ (7)
where, $s_{t}$ represents the state at time step $t$; $a_{t}$ represents the action at time step $t$; $r_{t+1}$ represents the immediate reward obtained after taking action $a_{t}$; $T$ represents the length of the trajectory; $s_{0}$ is the initial state, and $s_{T}$ is the terminal state. The cumulative reward $R(\tau)$ of a trajectory $\tau$ is obtained by summing the rewards $r_{t}$ obtained at each time step $t$:
$R(\tau)=\sum_{t=0}^{T}r_{t},$ (8)
The objective function $J(\pi_{\theta})$ in the PPO algorithm aims to maximize the expected cumulative reward obtained by following the policy $\pi_{\theta}$. Mathematically, this objective function is expressed as:
$J(\pi_{\theta})=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)],$ (9)
where $J(\pi_{\theta})$ represents the expected return under policy $\pi_{\theta}$, with the expectation taken over trajectories sampled from the policy distribution $\pi_{\theta}$. The term $R(\tau)$ denotes the total reward accumulated along a trajectory $\tau$. The policy gradient $\nabla_{\theta}J(\pi_{\theta})$, indicating the direction of steepest ascent in the objective function $J(\pi_{\theta})$, is computed using the policy gradient theorem. The gradient of the expected return is expressed as the expectation of the gradient of the log-probability of trajectories weighted by the return:
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\tau\sim\pi_{\theta}}[\nabla_{\theta}\log P(\tau\mid\theta)R(\tau)],$ (10)
where $\nabla_{\theta}\log P(\tau\mid\theta)$ represents the gradient of the log-probability of trajectories with respect to the policy parameters $\theta$. This computation involves sampling trajectories from the policy distribution $\pi_{\theta}$ and weighting the log-probability gradients by their respective returns. The policy parameters $\theta$ are updated iteratively using gradient ascent, with the update rule:
$\theta_{k+1}=\theta_{k}+\alpha\nabla_{\theta}J(\pi_{\theta}),$ (11)
where $\theta_{k}$ represents the policy parameters at the $k$-th iteration, $\alpha$ is the learning rate controlling the step size, and $\nabla_{\theta}J(\pi_{\theta})$ denotes the policy gradient.
In the PPO algorithm, the trajectories generated by the agent are instrumental in shaping the policy update process. The objective function, $J(\pi_{\theta})$, is designed to enhance the agent’s performance by maximizing the expected cumulative reward under the policy $\pi_{\theta}$. This optimization involves iteratively updating the policy parameters $\theta$ using the policy gradient $\nabla_{\theta}J(\pi_{\theta})$, while maintaining stability and preventing substantial policy shifts. The policy gradient is estimated using a collection of trajectories, $\mathcal{D}=\{\tau_{i}\}_{i=1,\ldots,N}$, each generated by the agent interacting with the environment according to the policy. The policy gradient can be calculated as follows:
$\hat{g}=\frac{1}{|\mathcal{D}|}\sum_{\tau\in\mathcal{D}}\sum_{t=0}^{T}\nabla_{\theta}\log\left(\pi_{\theta}(a_{t}\mid s_{t})\right)R(\tau),$ (12)
where $|\mathcal{D}|$ represents the number of trajectories in $\mathcal{D}$, and $T$ is the length of each trajectory. This formulation offers a simplified approach to computing the policy gradient. Moreover, PPO commonly employs a clipped surrogate objective function to facilitate more stable policy updates. This objective is expressed as:
$L^{CLIP}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{\tau\in\mathcal{D}}\sum_{t=0}^{T}\min\left(r_{t}(\theta)A_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right),$ (13)
where $r_{t}(\theta)$ is the likelihood ratio between the old and new policies, $A_{t}$ denotes the advantage function, and $\epsilon$ is a hyperparameter that defines the clipping range. The inclusion of this clipped surrogate objective aids in moderating the extent of policy updates, thereby stabilizing the training process.
In summary, the PPO algorithm presents a robust framework for the training of DRL agents. It operates through an iterative process of updating policy parameters aimed at maximizing the expected cumulative reward. Crucially, PPO maintains stability and mitigates the risk of substantial policy deviations. Employing policy gradients along with clipped surrogate objectives, PPO ensures effective and stable training across a diverse array of reinforcement learning tasks. This methodology not only enhances the efficiency of learning but also contributes significantly to the consistency and reliability of the training process.
@@ II.4 DRL-Enhanced Active Flow Control
This section outlines the design of a framework utilizing DRL algorithms to address AFC problems. Firstly, it is necessary to define two fundamental elements in DRL: the agent and the environment. The agent can be designed as algorithms suitable for solving continuous action control problems, such as PPO agent or SAC agent. The CFD simulation of the flow around an elliptical cylinder is chosen as the environment for the agent to interact with. Rabault et al.Rabault et al. (2020) presented the main structure of the custom environment class, encouraging readers to delve into the detailed content. This framework design applies DRL methods to solve the AFC problem, enabling the agent to continuously learn and optimize its behavior through interaction with the environment to achieve the control goal. Fig. 4 illustrates the fundamental framework of DRL, the computational process of a single episode during training, and the formulas utilized.
Figure 4: DRL leverages the concurrent execution of multiple environments within a parallel computational framework. Each iteration consists of two distinct phases: exploration and optimization. During the exploration phase, the agent engages with the environment via its current policy $\pi$, executing actions $a_{t}$, receiving states $s_{t}$ and rewards $r_{t}$, and noting terminal states. Upon the result of the iteration, the parallel computation across $n$ environments yields a set of trajectories $\{\tau_{i}\}_{i=1}^{n}$, with each trajectory $\tau_{i}$ comprising the sequence of states, actions, and rewards as $\tau_{i}=\{(s_{0},a_{0}),(s_{1},a_{1},r_{1}),\ldots,(s_{T},a_{T},r_{T})\}$, where $(s_{t},a_{t},r_{t})$ represents the state, action , and reward at time-step $t$, respectively, for the $i$-th environment. In the optimization phase, the agent samples a minibatch of trajectories from the experience replay buffer to compute the policy gradient $\nabla_{\theta}J(\pi_{\theta})$. The parameters $\theta$ of the policy network are updated using gradient ascent, i.e., $\theta\leftarrow\theta+\alpha\nabla_{\theta}J(\pi_{\theta})$, where $\alpha$ denotes the learning rate. This process iterates, with the agent deploying the updated policy $\pi_{\theta^{\prime}}$ to interact with the environment, computing the advantage function $A(s_{t},a_{t})$, and recalculating the policy gradient for subsequent updates of $\theta$. Through iterative refinement, the policy network incrementally enhances its action selection strategy within the environment.
  • State $s_{t}$: Traditional state spaces are often too expansive for effective policy learning, requiring operations in partially observable environments. As a result, an agent’s performance heavily relies on the quality and relevance of the data it can observe. The impact of partial observability of states $s_{t}$ on control results was mentioned in Viquerat et al.Viquerat et al. (2022)’s research.Integrating instantaneous velocity or pressure information from CFD simulations to characterize the environmental state $s_{t}$, the DRL agent can effectively perceive and analyze the current flow conditions. To achieve the control objectives of reducing drag and lift around the elliptical cylinder and controlling the shedding of vortices in the wake field, multiple probes are strategically placed around the elliptical cylinder and within its wake field. This method is advantageous for gathering comprehensive data pertaining to the flow dynamics around the cylinder and subsequent wake characteristics. The design of probes placement in this study was based on the design principles articulated by Rabault et al.Rabault et al. (2019), which advocate strategic positioning of probes around the cylinder and within the wake region of the cylinder.In their studies, Wang and XuWang and Xu (2024) and Wang et al.Wang et al. (2022a) all referred to the probes arrangement scheme proposed by Rabault et al.Rabault et al. (2019).
@@
  • Action $a_{t}$ : The control action involves the blowing and suction of two synthetic jets on the elliptical cylinder. The flow rates of the two synthetic jets are constrained such that $Q_{\text{lower}}+Q_{\text{upper}}=0$. Therefore, the RL agent manipulates the velocity of the upper synthetic jets, while the lower synthetic jets has the same velocity but in the opposite direction. The velocity of the synthetic jets is set to be less than 2% of the inlet mean velocity. This restriction is imposed to prevent the agent from taking shortcuts and using excessively high jet velocities to achieve drag reduction, which contradicts the fundamental principle of AFC: energy efficiency. Furthermore, each control action lasts for 50 time steps, termed as a viscous action. Allowing faster actions could harm overall performance as they need time to influence the environment. Ensuring smooth transitions between current and updated actions is crucial. Let $a_{t}$ and $a_{t+1}$ represent the action magnitudes at consecutive time steps. To ensure a gradual transition between these actions and mitigate abrupt changes in jet velocity, a smoothing function $S$ can be defined mathematically. The function $S$ interpolates between the previous action magnitude $V_{\Gamma_{1},T_{i-1}}$ and the target action magnitude $a$ at time step $i$, resulting in an updated value $V_{\Gamma_{1},T_{i}}$. This smoothing function can be expressed as:
@@
  • Reward $r_{t}$ : This study incorporates the average drag coefficient (without control), the time-averaged value of the drag coefficient, and the time-averaged value of the lift coefficient (averaged over the duration of the action). The primary objective is to minimize the drag while maintaining a low lift coefficient; hence, the weight of the drag coefficient in the reward function is set to 1, while the lift coefficient is assigned a weight $w$, typically chosen to be 0.1 or 0.2. This hyperparameter $w$ is determined through tuning. Including a constant (the average drag coefficient without control) in the reward function helps guide it towards positive rewards, which positively affects the convergence speed of the learning curve. The reward function $f(\cdot)$ translates the drag and lift coefficients into a numerical reward assessing the agent’s behavior. The specific reward function takes the following form:
@@
    $C_{D,0}$ represents the baseline drag coefficient, serving as a reference point for drag reduction. $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective of minimizing this value relative to the baseline. $(C_{L})_{T_{i}}$ signifies the lift coefficient at time step $T_{i}$, whose absolute value is penalized to mitigate lift forces that may destabilize the flow around the bluff body. $\omega$ is a weighting factor that quantifies the trade-off between minimizing drag and controlling lift fluctuations.
  • Parameter. Time granularity in the DRL framework is crucially set based on the fluid dynamics problems studied and significantly affects agent performance and learning task complexity. Precise adjustments of this hyperparameter, informed by prior research and our ongoing strategy refinement, are vital for achieving optimized control outcomes. Each action is applied for a duration of 0.025 seconds ( Corresponds to 25 numerical simulation time steps ), and each episode consists of 100 control actions. The maximum total time for an episode is defined as $T_{\max}=T_{100}=2.5$ nondimensional time units. The shedding frequency of vortices around the elliptical cylinder in this study ranges from 0.3 to 0.4, corresponding to 6.25 to 8.33 vortex shedding cycles per episode. The length of episodes is intentionally chosen to encompass multiple shedding cycles, facilitating effective learning of the control algorithm by the agent through observation and adaptation of flow dynamics over several cycles.
II.5 High-Performance Computing in DRL and CFD
The DRL training and numerical simulation in this study were conducted on a high-performance computing (HPC) system equipped with an Intel® Xeon® Platinum 8358 CPU, operating at 2.60 GHz and comprising a total of 64 cores distributed evenly across two sockets, each containing 32 cores. Numerical simulations were performed using the eighth version of the open-source software platform, OpenFOAM® version 8.Jasak et al. (2007) The deployment of DRL algorithms was based on the Tensorforce library.Schaarschmidt et al. (2018) The ANN definitions and gradient descent algorithms were facilitated using the TensorFlow open-source library.Abadi et al. (2016) For DRL implementation, agent was loaded using the Tensorforce platform with the PPO algorithm, and the environment was constructed using the Gym interface.Brockman et al. (2016) Fig. 4 illustrates the use of CFD calculations as the environment for DRL training and shows the effects of using 60 parallel CFD environments to accelerate the training process. This parallel strategy was inspired by the research of Rabault and KuhnleRabault and Kuhnle (2019a) and Wang et al.Wang et al. (2022a), and also incorporated insights from the work of Wang et al.Wang et al. (2022a) For a more comprehensive understanding, readers are encouraged to consult their works and the references therein.

### v3

@@ II.3 Deep reinforcement learning
The structure of our employed DRL model is illustrated in Fig. 3. the DRL marries the perceptual power of deep learning with the decision-making and optimization capabilities of reinforcement learning, effectively overcoming their individual limitations and enhancing overall performance. Reinforcement Learning serves as the core mechanism for facilitating learning through the interaction between an agent and its environment. At each time step $t$, the agent observes the current state $s_{t}$ of the environment, which contains all the necessary information for decision-making. Based on this observation, the agent selects and executes an action $a_{t}$ according to its policy $\pi$. The policy, serving as a strategy mapping states to actions, can be deterministic or probabilistic in nature. After the action is executed, the environment transitions to a new state $s_{t+1}$ and provides the agent with an immediate reward $r_{t+1}$. This reward, expressed as a numerical value, assesses the effectiveness of the action in progressing towards the agent’s objectives. Consequently, the agent incorporates this feedback-comprising both the reward and the new state—to refine its policy, utilizing Reinforcement Learning algorithms designed to optimize the accumulation of long-term rewards.
Figure 3: DRL combines deep learning’s capability to interpret complex data with reinforcement learning’s strategy to make informed decisions.
Proximal policy optimization
PPO algorithm algorithm is an improvement over the policy gradient methods. The core idea of the PPO algorithm is to update the policy through two crucial steps: data sampling and policy optimization. In the data sampling phase, we interact with the environment using the current policy and collect a batch of trajectory data denoted as $\tau$, which can be mathematically expressed as:
$\tau=(s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},s_{2},\ldots,s_{T-1},a_{T-1},r_{T},s_{T}).$ (11)
Here, $s_{t}$ represents the state at time step $t$, $a_{t}$ represents the action taken at time step $t$, and $r_{t}$ represents the reward received at time step $t$. The trajectory $\tau$ consists of a sequence of state-action-reward tuples observed during the interaction between the agent and the environment over a certain time horizon $T$. For a given trajectory $\tau$, at each interaction between the agent and the environment, the agent receives a reward value $r_{t}$ at time step $t$. Hence, the cumulative reward $R(\tau)$ of the trajectory $\tau$ is obtained by summing up the rewards obtained at each time step:
$R(\tau)=\sum_{t=0}^{T}r_{t}.$ (12)
Here, $T$ represents the time horizon or the length of the trajectory, and $r_{t}$ represents the reward received at time step $t$. The cumulative reward $R(\tau)$ provides a measure of the total reward obtained by the agent throughout the trajectory $\tau$.
In the policy optimization phase, we use the sampled data obtained in the previous phase to update the policy. The method used to update the policy is gradient ascent. To perform gradient ascent, we compute the gradient of the PPO objective function with respect to the policy parameters. The objective function of the PPO algorithm is a surrogate objective designed to strike a balance between exploration and exploitation. It can be expressed as follows:
$\left.L^{CLIP}(\theta)=\hat{\mathbb{E}}_{t}\left[\min\left(r_{t}(\theta)\right)\hat{A}_{t},\operatorname{clip}\left(r_{t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{t}\right)\right],$ (13)
where $\theta$ represents the policy parameters that are being optimized during the policy update process. $\hat{\mathbb{E}}_{t}$ denotes the empirical expectation, indicating that the expectation is estimated using the samples collected from the environment. $r_{t}(\theta)$ represents the ratio of the probabilities between the new policy and the old policy, evaluated at time step $t$. $\hat{A}_{t}$ corresponds to the advantage estimate at time step $t$. The advantage represents the advantage of taking a particular action compared to the expected value of the action under the current policy. The $\operatorname{clip}(.)$ function performs clipping on the policy ratio to ensure that the update does not deviate too much from the old policy. It restricts the ratio within the range $[1-\varepsilon,1+\varepsilon]$, where $\varepsilon$ is a small value. $[\min\left(r_{t}(\theta)\right)\hat{A}_{t},\operatorname{clip}\left(r_{t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{t}]$ expression represents the clipped surrogate objective. It consists of two terms: the first term is the minimum between $r_{t}(\theta)\hat{A}_{t}$ and $\operatorname{clip}\left(r_{t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{t}$, and the second term is the clipped version of $r_{t}(\theta)\hat{A}_{t}$. The purpose of this formulation is to balance policy improvement while ensuring a controlled update magnitude.
@@ II.4 DRL-enhanced active flow control
Our AFC framework utilizes DRL algorithms, consisting of the agent and the environment. The agent, designed for continuous action control, employs algorithms like PPO or SAC. The environment chosen for interaction with the agent is the CFD simulation of the flow around an elliptical cylinder. The custom environment class, as presented by Rabault et al.Rabault et al. (2020), provides the main structure, and further details can be explored in their work. Figure 4 illustrates the fundamental framework of DRL, including the computational process of a single training episode and the utilized formulas.
Figure 4: The Proximal Policy Optimization algorithm is trained using a parallel computing framework to simultaneously run multiple environments. In the interaction stage between agent and environment, an agent interacts with its environments by following the current policy $\pi$, where it executes actions $a_{t}$, receives states $s_{t}$, and gathers rewards $r_{t}$, identifying terminal states in the process. This interaction generates a set of trajectories $\{\tau_{i}\}_{i=1}^{n}$, each $\tau_{i}$ being a sequence of state-action-reward tuples ending at a terminal state $T$. The agent uses these trajectories to compute the policy gradient $\nabla_{\theta}J(\pi_{\theta})$ and applies gradient ascent to update the policy parameters $\theta$, aiming to maximize cumulative rewards. Through repeated iterations, leveraging updated policies and recalculated gradients, the agent continuously enhances its policy effectiveness, optimizing its performance across various environments.
Fundamental components configuration
In DRL-Enhanced active flow control, it involves state ($s_{t}$), action ($a_{t}$), reward ($r_{t}$), and parameter, which are described as below. The traditional state space $s_{t}$ poses challenges in learning policies due to its large size, emphasizing the reliance of the agent’s performance in partially observable environments on the quality and relevance of observed data, as highlighted by Viquerat et al.Viquerat et al. (2022). To achieve the control objectives of reducing drag and lift and controlling vortex shedding in the wake field of the elliptical cylinder, strategically placed probes are utilized. The design of probe placement follows the principles proposed by Rabault et al.Rabault et al. (2019). This scheme was also adopted by Wang and XuWang and Xu (2024a) and Wang et al.Wang et al. (2022a) in their respective studies.
@@
The control action $a_{t}$ involves blowing and suction of two synthetic jets on the elliptical cylinder. The flow rates of the upper and lower synthetic jets are constrained such that their sum is zero, i.e., $Q_{\text{lower}}+Q_{\text{upper}}=0$. The RL agent controls the velocity of the upper synthetic jet, while the lower synthetic jet has the same velocity but in the opposite direction. To ensure energy efficiency in AFC, the velocity of the synthetic jets is limited to less than 2% of the inlet mean velocity, preventing the agent from utilizing excessively high jet velocities for drag reduction. Each control action, known as a ‘viscous action’, has a duration of 50 time steps. To ensure smooth transitions between consecutive actions, a smoothing function $S$ is introduced:
@@
The reward $r_{t}$ evaluates the agent based on the average drag coefficient, time-averaged drag coefficient, and time-averaged lift coefficient during the action, with the objective of minimizing drag while keeping the lift coefficient low. The weight of the drag coefficient in the reward function is set to 1, indicating its importance, while the lift coefficient is assigned a weight $w$. The specific reward function takes the following form:
@@
where $C_{D,0}$ represents the baseline drag coefficient, serving as a reference point for drag reduction. $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective of minimizing this value relative to the baseline. $(C_{L})_{T_{i}}$ signifies the lift coefficient at time step $T_{i}$, whose absolute value is penalized to mitigate lift forces that may destabilize the flow around the bluff body. $\omega$ is a weighting factor that quantifies the trade-off between minimizing drag and controlling lift fluctuations.
The time granularity parameter significantly affects the agent’s performance and the complexity of the learning task. Precise adjustment to this hyperparameter, based on prior research and ongoing strategy refinement, is essential for achieving optimized control outcomes. In particular, each action in our analysis is applied for a duration of 0.025 seconds, which corresponds to 25 numerical simulation time steps. Each episode consists of 100 control actions, resulting in a maximum total time for an episode of $T_{\max}=T_{100}=2.5$ nondimensional time units. The shedding frequency of vortices around the elliptical cylinder in this study ranges from 0.3 to 0.4, which corresponds to 6.25 to 8.33 vortex shedding cycles per episode. The length of episodes is intentionally chosen to encompass multiple shedding cycles, which facilitates effective learning of the control algorithm by allowing the agent to observe and adapt to the flow dynamics over several cycles.
Computing specifications
The DRL training and numerical simulation are conducted on a high-performance computing (HPC) system equipped with an Intel® Xeon® Platinum 8358 CPU, operating at 2.60 GHz, and comprising a total of 64 cores evenly distributed across two sockets, each containing 32 cores. Specifically, the numerical simulation is performed using version 8 of the open-source software platform OpenFOAM® version 8 Jasak et al. (2007). The deployment of DRL algorithms is based on theTensorforce library Schaarschmidt et al. (2018), while the definitions of artificial neural networks (ANNs) and gradient descent algorithms are facilitated using the TensorFlow open-source library Abadi et al. (2016). For the DRL implementation, the agent is loaded using the Tensorforce platform with the PPO algorithm, and the environment is constructed using the Gym interface Brockman et al. (2016) . Fig. 4 illustrates the use of CFD calculations as the environment for DRL training and shows the effects of using 60 parallel CFD environments to accelerate the training process. This parallel strategy is inspired by the studies of Rabault and KuhnleRabault and Kuhnle (2019a) and Wang et al.Wang et al. (2022a), and also incorporateds insights from the work of Wang et al.Wang et al. (2022a) For a more comprehensive understanding, readers are encouraged to consult their works and the references therein.